# GI-NAS: Boosting Gradient Inversion Attacks through Adaptive Neural Architecture Search

### v2

@@ V-D Further Analysis
Dataset Model Random Seed Initial Gradient Matching Loss (×$\times 10^{-3}$) $\sigma_{loss}$ (×$\times 10^{-4}$) PSNR $\sigma_{psnr}$ CIFAR-10 $G_{c4}$ $s_{c4}$ 14.5338 0.0018 26.4506 0.1925 $s_{c5}$ 14.5335 26.2082 $s_{c6}$ 14.5336 26.0703 $G_{c5}$ $s_{c4}$ 13.4751 0.0018 30.1240 0.2675 $s_{c5}$ 13.4753 30.2045 $s_{c6}$ 13.4755 30.6222 $G_{c6}$ $s_{c4}$ 13.2704 0.0015 31.3139 0.0566 $s_{c5}$ 13.2701 31.4256 $s_{c6}$ 13.2702 31.3540 ImageNet $G_{i4}$ $s_{i4}$ 23.1931 4.6280 21.8376 0.1164 $s_{i5}$ 23.6335 22.0556 $s_{i6}$ 22.7083 21.8760 $G_{i5}$ $s_{i4}$ 19.5264 1.1102 23.6033 0.1889 $s_{i5}$ 19.5677 23.2439 $s_{i6}$ 19.7360 23.3226 $G_{i6}$ $s_{i4}$ 25.9479 3.6054 20.4730 0.1655 $s_{i5}$ 25.3497 20.8007 $s_{i6}$ 25.9975 20.6776
@@
TABLE X: Averaged architectural statistics of the searched optimal models by our NAS strategy under various settings tested in Section V-C. The default settings are highlighted in gray color. $\Delta_{connection}$ and $\Delta_{upsampling}$ are respectively the gains of the number of skip connections and the kernel size of upsampling modules when compared with the default settings.
@@
Preferences of NAS Outcomes. To figure out the implications behind NAS preferences, we trace back the searched optimal models from Section V-C and analyze their averaged architectural statistics in Table X. Note that the bicubic operation takes more neighbourhood pixels than the bilinear operation ($16>4$) and is thus considered to be more complex. Besides, the pixel shuffle operation rearranges lots of pixels and is also considered to be more complicated than the bilinear operation. From Table X, we observe that the optimal models are largely dependent on recovery difficulties. Generally, batches with larger sizes or under defenses are more difficult to recover. Compared to the default settings (highlighted in gray color) that are with the batch size $B=4$ and under no defenses, the more difficult settings would prefer models with higher complexity (e.g., more skip connections, larger kernel sizes, more complex operations). The complexity gains $\Delta_{connection}$ and $\Delta_{upsampling}$ are often greater than 00 when the settings are $B>4$ or under defenses. Easier batches prefer simpler, lighter architectures and harder batches prefer more complex, heavier architectures. This might be used to further improve the search efficiency. By considering the attack difficulties ahead of time, we can have a better estimation of the complexity of optimal models. Thus, we only need to focus on models with a certain type of complexity and skip other unrelated models even when there is no enough time for NAS.
Ablation Study. In Table XI, We report the PSNR results of GION and different variants of GI-NAS. GI-NAS optimizes on a fixed over-parameterized network with the search size $n=1$, which means that it is essentially the same as GION. GI-NAS only searches the skip connection patterns while GI-NAS only searches the upsampling modules. We observe that the performance of GI-NAS is very close to that of GION, as both of them adopt a changeless architecture. GI-NAS and GI-NAS indeed improve the attacks on the basis of GION or GI-NAS, which validates the contributions of individual search types. GI-NAS combines the above two search types and thus performs the best among all the variants.
TABLE XI: Ablation Study. We report the PSNR results of GION and different variants of GI-NAS on CIFAR-10 ($32\times 32$) and ImageNet ($256\times 256$) with the default batch size $B=4$.
Generalizability to More FL Global Models. We show the performance of different methods when attacking more FL global models (e.g., LeNet-Zhu [12], ConvNet-32 [14]) in Figure 6. We note that GI-NAS realizes the best reconstruction results in all the tested cases. These results further demonstrate the reliability of GI-NAS. We also find that the attack results on different FL global models are distinctively dissimilar even when applying the same gradient inversion method. This implies that future study may have a deeper look into the correlation between the gradient inversion robustness and the architecture of FL global model, and thus design more securing defense strategies from the perspective of adaptive network architecture choice.
(a) CIFAR-10
(b) ImageNet
Figure 6: PSNR results of GI-NAS and state-of-the-art gradient inversion methods when attacking more FL global models on CIFAR-10 ($32\times 32$) and ImageNet ($256\times 256$) with the default batch size $B=4$.
@@ VI Conclusion
In this paper, we propose GI-NAS, a novel gradient inversion method that makes deeper use of the implicit architectural priors for gradient inversion. We first systematically analyze existing gradient inversion methods and emphasize the necessity of adaptive architecture search. We then build up our model search space by designing different upsampling modules and skip connection patterns. To reduce the computational overhead, we leverage the initial gradient matching loss as the training-free search metric to select the optimal model architecture and provide extensive experimental evidence that such a metric highly correlates with the real attack performance. Extensive experiments prove that GI-NAS can achieve state-of-the-art performance compared to existing methods, even under more practical FL scenarios with high-resolution images, large-sized batches, and advanced defense strategies. We also provide deeper analysis on various aspects, including time costs, robustness to network parameters initialization, robustness to latent codes initialization, implications behind searched results, ablation study on each search type, and generalizability to more FL global models. We hope that the remarkable attack improvements of GI-NAS over existing gradient inversion methods may help raise the public awareness of such privacy threats, as the sensitive data would be more likely to be revealed or even abused. Moreover, we hope that the idea of this paper may shed new light on the gradient inversion community and facilitate the research in this field.
@@
Limitations and Future Directions. Although we have empirically demonstrated that the initial gradient matching loss is highly negatively correlated to the final attack performance, it is still of great significance to further prove the effectiveness of this search metric with rigorous theoretical analysis. In the future work, we hope to provide more insightful theoretical results with regard to this search metric from various perspectives, such as the frequency spectrum [86, 74] or the implicit neural architectural priors [22, 87].

### v3

@@ V-D Further Analysis
Dataset Model Random Seed Initial Gradient Matching Loss (×$\times 10^{-3}$) $\sigma_{loss}$ (×$\times 10^{-4}$) PSNR $\sigma_{psnr}$ CIFAR-10 $G_{C1}$ $Seed_{C1}$ 8.5647 0.0012 28.7283 0.1211 $Seed_{C2}$ 8.5648 28.8082 $Seed_{C3}$ 8.5645 28.5703 $G_{C2}$ $Seed_{C1}$ 8.4321 0.0006 33.5848 0.2581 $Seed_{C2}$ 8.4323 33.4631 $Seed_{C3}$ 8.4322 33.9584 $G_{C3}$ $Seed_{C1}$ 7.9501 0.0015 49.4290 0.2936 $Seed_{C2}$ 7.9499 49.1171 $Seed_{C3}$ 7.9502 49.7039 ImageNet $G_{I1}$ $Seed_{I1}$ 21.0167 5.6866 22.9013 0.0329 $Seed_{I2}$ 22.1401 22.8912 $Seed_{I3}$ 21.4250 22.8398 $G_{I2}$ $Seed_{I1}$ 24.4600 2.2089 21.1475 0.3040 $Seed_{I2}$ 24.0722 21.6449 $Seed_{I3}$ 24.0828 21.6989 $G_{I3}$ $Seed_{I1}$ 18.5319 6.0290 23.5278 0.1884 $Seed_{I2}$ 18.3819 23.2974 $Seed_{I3}$ 17.4208 23.6709
Robustness to Network Parameters Initialization. Following the general paradigm of previous training-free NAS methods [61, 62, 63] in other areas, the parameters of all the NAS candidates are randomly initialized and there is no special need to tune the parameters before calculating the initial gradient matching losses. Moreover, GI-NAS records both the architectures as well as the initial parameters when finding the optimal models. Therefore, for the model we decide on, its initial gradient matching loss in the NAS time (i.e., Stage 1 as shown in Figure 2) is the same as its initial gradient matching loss in the optimization time (i.e., Stage 2 as shown in Figure 2), and the influences of parameters initialization can be further reduced. To further demonstrate the robustness, we randomly select models with different architectures and apply different parameter initialization seeds to them to attack the same batch. From Table XI, we find that for the same architecture, both the initial gradient matching losses and the final PSNR results stay within a steady range when changing the parameter initialization seeds. However, when the architectures change, both the initial gradient matching losses and the final PSNR results will vary significantly. The impacts of different architectures are much greater than the impacts of different parameter initialization seeds. As a result, we conclude that GI-NAS is stable and insensitive to network parameters initialization. We also observe that the models with smaller initial gradient matching losses often have better final PSNR results in Table XI, which is consistent with our aforementioned findings in Figure 5 and Table VI. This again validates the rationality and usefulness of adopting the initial gradient matching loss as the training-free search metric.
Robustness to Latent Codes Initialization. The latent code $\mathbf{z_{0}}$ is randomly sampled from the Gaussian distribution $\mathcal{N}(0,1)$ and then kept frozen. Thus, we explore the influences of different latent codes by changing their initialization seeds. From Table XII, we discover that although different latent code initialization seeds on the same architecture may result in different initial gradient matching losses, the differences are still very minor (less than $10^{-6}$ for CIFAR-10 and less than $10^{-3}$ for ImageNet), and the impacts on the final PSNR results are limited (less than $0.5$ dB for CIFAR-10 and less than $0.4$ dB for ImageNet). However, when using different architectures, both these two values will change dramatically, and the differences are much greater than those brought by changing the latent code initialization seeds. Hence, GI-NAS is also robust to latent codes initialization.
TABLE XII: Results when attacking the same batch by different networks using different latent code initialization seeds on CIFAR-10 ($32\times 32$) and ImageNet ($256\times 256$) with the default batch size $B=4$. The parameter initialization seeds are fixed for all the models. $G_{C4}\sim G_{C6}$ and $G_{I4}\sim G_{I6}$ are randomly sampled from the model search space, while $Seed_{C4}\sim Seed_{C6}$ and $Seed_{I4}\sim Seed_{I6}$ are randomly selected and are different from each other. Note that $\sigma_{loss}$ and $\sigma_{psnr}$ are respectively the standard deviations of the initial gradient matching losses and PSNR results on each network.
Dataset Model Random Seed Initial Gradient Matching Loss (×$\times 10^{-3}$) $\sigma_{loss}$ (×$\times 10^{-4}$) PSNR $\sigma_{psnr}$ CIFAR-10 $G_{C4}$ $Seed_{C4}$ 14.5338 0.0018 26.4506 0.1925 $Seed_{C5}$ 14.5335 26.2082 $Seed_{C6}$ 14.5336 26.0703 $G_{C5}$ $Seed_{C4}$ 13.4751 0.0018 30.1240 0.2675 $Seed_{C5}$ 13.4753 30.2045 $Seed_{C6}$ 13.4755 30.6222 $G_{C6}$ $Seed_{C4}$ 13.2704 0.0015 31.3139 0.0566 $Seed_{C5}$ 13.2701 31.4256 $Seed_{C6}$ 13.2702 31.3540 ImageNet $G_{I4}$ $Seed_{I4}$ 23.1931 4.6280 21.8376 0.1164 $Seed_{I5}$ 23.6335 22.0556 $Seed_{I6}$ 22.7083 21.8760 $G_{I5}$ $Seed_{I4}$ 19.5264 1.1102 23.6033 0.1889 $Seed_{I5}$ 19.5677 23.2439 $Seed_{I6}$ 19.7360 23.3226 $G_{I6}$ $Seed_{I4}$ 25.9479 3.6054 20.4730 0.1655 $Seed_{I5}$ 25.3497 20.8007 $Seed_{I6}$ 25.9975 20.6776
@@
TABLE XIII: Averaged architectural statistics of the searched optimal models by our NAS strategy under various settings tested in Section V-C. The default settings are highlighted in gray color. $\Delta_{connection}$ and $\Delta_{upsampling}$ are respectively the gains of the number of skip connections and the kernel size of upsampling modules when compared with the default settings.
@@
Preferences of NAS Outcomes. To figure out the implications behind NAS preferences, we trace back the searched optimal models from Section V-C and analyze their averaged architectural statistics in Table XIII. Note that the bicubic operation takes more neighbourhood pixels than the bilinear operation ($16>4$) and is thus considered to be more complex. Besides, the pixel shuffle operation rearranges lots of pixels and is also considered to be more complicated than the bilinear operation. From Table XIII, we observe that the optimal models are largely dependent on recovery difficulties. Generally, batches with larger sizes or under defenses are more difficult to recover. Compared to the default settings (highlighted in gray color) that are with the batch size $B=4$ and under no defenses, the more difficult settings would prefer models with higher complexity (e.g., more skip connections, larger kernel sizes, more complex operations). The complexity gains $\Delta_{connection}$ and $\Delta_{upsampling}$ are often greater than 00 when the settings are $B>4$ or under defenses. Easier batches prefer simpler, lighter architectures and harder batches prefer more complex, heavier architectures. This might be used to further improve the efficiency. By considering the difficulties ahead of time, we can have a better estimation of the complexity of optimal models. Thus, we only need to focus on models with a certain type of complexity and skip other unrelated models even when there is not enough time for NAS.
@@ VI Conclusion
In this paper, we propose GI-NAS, a novel gradient inversion method that makes deeper use of the implicit architectural priors for gradient inversion. We first systematically analyze existing gradient inversion methods and emphasize the necessity of adaptive architecture search. We then build up our model search space by designing different upsampling modules and skip connection patterns. To reduce the computational overhead, we leverage the initial gradient matching loss as the training-free search metric to select the optimal model architecture and provide extensive experimental evidence that such a metric highly correlates with the real attack performance. Extensive experiments show that GI-NAS can achieve state-of-the-art performance compared to existing methods, even under more practical FL scenarios with high-resolution images, large-sized batches, and advanced defense strategies. We also provide deeper analysis on various aspects, such as time costs, ablation studies, generalizability to more FL global models, robustness to network parameters initialization, robustness to latent codes initialization, and implications behind the NAS searched results. We hope that the remarkable attack improvements of GI-NAS over existing gradient inversion methods may help raise the public awareness of such privacy threats, as the sensitive data would be more likely to be revealed or even abused. Moreover, we hope that the idea of this paper may shed new light on the gradient inversion community and facilitate the research in this field.
@@
Limitations and Future Directions. Although we have empirically demonstrated that the initial gradient matching loss is highly negatively correlated to the final attack performance, it is still of great significance to further prove the effectiveness of this search metric with rigorous theoretical analysis. In the future work, we hope to provide more insightful theoretical results with regard to this search metric from various perspectives, such as the frequency spectrum [77, 67] or the implicit neural architectural priors [21, 78]. Furthermore, inspired by the effectiveness of NAS for attacks in this paper, a promising defensive direction involves exploring victim-side neural architecture search to discover model structures inherently more resistant to gradient inversion while maintaining model accuracy, or developing adaptive gradient perturbation techniques on the FL global model guided by NAS principles.
Appendix A Attack Results on More Datasets
We further compare the attack effectiveness of GI-NAS with other methods on more datasets, including FFHQ [79], MNIST [80], and SVHN [81]. These datasets are also considered in previous gradient inversion methods [82, 12]. From Table XIV, we observe that GI-NAS consistently outperforms all the compared methods across diverse datasets. This consistent superiority suggests that GI-NAS is a highly generalizable and effective method for privacy attacks in federated learning.
TABLE XIV: Quantitative comparison of GI-NAS to state-of-the-art gradient inversion methods on FFHQ ($64\times 64$), MNIST ($28\times 28$), and SVHN ($32\times 32$) with the default batch size $B=4$.
Dataset Metric IG GI GGL GIAS GIFD GION GI-NAS FFHQ PSNR $\uparrow$ 12.1571 12.4204 10.8660 13.2693 13.7942 31.7174 36.2522 SSIM $\uparrow$ 0.2226 0.2237 0.1958 0.3418 0.3609 0.9865 0.9954 FSIM $\uparrow$ 0.6212 0.6128 0.6099 0.6623 0.6763 0.9879 0.9979 LPIPS $\downarrow$ 0.7448 0.7516 0.6327 0.6160 0.6182 0.0212 0.0025 MNIST PSNR $\uparrow$ 15.1129 14.1863 9.3894 22.8035 28.8421 31.9038 37.4773 SSIM $\uparrow$ 0.3497 0.3220 0.0670 0.5962 0.7874 0.9935 0.9953 FSIM $\uparrow$ 0.5319 0.5115 0.3957 0.7520 0.8959 0.9982 0.9993 LPIPS $\downarrow$ 0.3898 0.4425 0.5735 0.2240 0.0923 0.0014 0.0005 SVHN PSNR $\uparrow$ 20.9779 20.9475 16.9157 21.0760 22.8974 44.8053 46.4780 SSIM $\uparrow$ 0.6702 0.7047 0.4427 0.6704 0.7894 0.9774 0.9870 FSIM $\uparrow$ 0.7993 0.8121 0.6535 0.8045 0.8531 0.9895 0.9941 LPIPS $\downarrow$ 0.5145 0.5247 0.6789 0.5148 0.4270 0.0045 0.0028
Appendix B Attack Results on Heterogeneous Data
FL systems would commonly face the problem of heterogeneous data [83]. Thus, to further validate the robustness of our method, we strictly align with the prior works [84, 85] to design the data heterogeneity experiments. To be specific, we randomly allocate the classes of datasets into $N_{client}$ clients with a probability $q$ of assigning a training example with a certain label to a particular client and a probability $\frac{1-q}{N_{client}-1}$ of assigning it to other clients. When $q=\frac{1}{N_{client}}$, the client’s local training data are independent and identically distributed (IID), otherwise the client’s local training data are Non-IID.
Figure 6: Attack performance of GI-NAS when changing the values of $q$ (the probability of assigning a training example with a certain label to a particular client) on ImageNet ($256\times 256$) with the default batch size $B=4$. Note that there are $N_{client}=10$ clients in total, which indicates that if $q=\frac{1}{10}$, the distribution is IID. These settings strictly adhere to the settings in previous works [84, 85] considering data heterogeneity.
We set $N_{client}=10$, which means that if $q=\frac{1}{10}$, the distribution is IID. We select $q=0.10,0.30,0.50,0.70,0.90,0.99$, and test the performance of GI-NAS under these settings. From Figure 6, we notice that the attack performance of GI-NAS stays within a steady range when changing the values of $q$. Therefore, we conclude that GI-NAS is also robust to the influences of data heterogeneity.
Appendix C More Visualized Results
We provide more visualized comparisons in Figure 7. We observe that all the other methods struggle to conduct the reconstruction attacks, while our proposed GI-NAS again realizes the best performance in terms of visual fidelity.
Original
IG
GI
GGL
GIAS
GIFD
GION
GI-NAS
Figure 7: More qualitative results of GI-NAS and state-of-the-art gradient inversion methods on ImageNet ($256\times 256$) with the larger batch size $B=32$.